### v1

# Application of the microscopic optical potential of chiral effective field theory in astrophysical neutron-capture reactions

Abstract
The microscopic global nucleon–nucleus optical potential proposed by Whitehead, Lim, and Holt (WLH) is a state-of-the-art potential developed within the framework of many-body perturbation theory using realistic nuclear interactions from chiral effective field theory. Given its potentially greater predictive power for reactions involving exotic isotopes, we apply it to the calculations of astrophysical neutron-capture reactions for the first time, which are particularly important to the nucleosynthesis of elements heavier than iron. It is found that this potential provides a good description of experimental known neutron-capture cross sections and Maxwellian-averaged cross sections. For unstable neutron-rich nuclei, we comprehensively calculate the neutron-capture reaction rates for all nuclei with $26\leq Z\leq 84$, located between the valley of stability and the neutron drip line, using the backward-forward Monte Carlo method with the $f_{rms}$ deviation as the $\chi^{2}$ estimator. The results reveal a noticeable separation in the uncertainty of rates around an isospin asymmetry of 0.28 under the constraint $f_{rms}\leq 1.56$. This highlights the critical role of isospin dependence in optical potentials and suggests that future developments of the WLH potential may pay special attention to the isospin dependence.
journal: Physics Letters B
1 Introduction
The elements heavier than iron in the Universe are mostly synthesized through neutron-capture reactions [1, 2]. The slow ($s$) [3, 4, 5, 6] and rapid ($r$) [7, 8, 9, 10] neutron-capture processes with their neutron-capture timescales slower or more rapid, respectively, than $\beta$ decay timescale of unstable nuclei are responsible for the vast majority of heavy element abundances. Between these two processes, other neutron-capture processes, such as the intermediate ($i$) [11, 12, 13] and neutron ($n$) [14, 15] processes, may also be indispensable for explaining certain specific features in elemental abundance patterns. However, a large number of neutron-capture rates involved in these nucleosynthesis processes are not known experimentally [16, 17, 18], and only theoretical predictions can fill the gaps. Theoretical neutron-capture rates are commonly treated within the framework of the Hauser-Feshbach statistical model [19, 20]. The optical potential, nuclear level density, and $\gamma$-ray strength function are crucial nuclear structure inputs to this model. Significant efforts have been devoted to improving the descriptions of the nuclear level densities, both from phenomenological [21, 22, 23, 24, 25] and microscopic [26, 27, 28] perspectives, and the strength functions through phenomenological [29, 30, 31, 32, 33] and microscopic [34, 35, 36] approaches.
For the optical potential, which is essential for calculating the neutron transmission coefficient, a large number of developed models are available [37]. These include phenomenological potentials [38], semi-microscopic potentials [39, 40, 41], dispersive potentials [42], as well as models with uncertainty quantification [43, 44, 45]. However, most of them either rely on fitting nuclear reaction data, introducing unreliability when extrapolated to exotic nuclei lacking experimental constraints, or have limited applicability ranges, preventing their global use in astrophysical neutron-capture reactions [46]. The microscopic optical potentials developed from chiral effective field theory is a new booming area [47, 48, 49, 50, 51]. Recently, Whitehead, Lim, and Holt (WLH) constructed the full microscopic global nucleon-nucleus optical potential based on an analysis of 1800 isotopes within the framework of many-body perturbation theory using realistic nuclear interactions from chiral effective field theory [52]. This state-of-the-art potential, derived from the chiral nuclear force without any parameters fitted to nucleon-nucleus scattering data, is expected to be more powerful in predicting neutron-capture rates for nuclei far from stability. Moreover, it provides quantified uncertainties for the parameters, enabling one to estimate the uncertainty in a given reaction observable. While the WLH potential displays impressive capability in describing elastic scattering [52], charge-exchange reactions [53], and transfer reactions [54], it has not yet been widely applied to astrophysical neutron-capture reactions.
In this work, we apply the WLH potential to the study of astrophysical neutron-capture reactions for the first time and test its applicability to nuclei with $26\leq Z\leq 84$ lying between the neutron drip line and the valley of $\beta$ stability, which is a critical region for the nucleosynthesis of elements heavier than iron. The rest of this article is organized as follows. In Sec. 2, the nuclear theoretical models used in our work are described. In Sec. 3, the calculated results are compared with experimental data, and the WLH potential is applied to neutron-rich nuclei. In Sec. 4, the conclusions are presented.
2 Theoretical model
We use the Hauser-Feshbach statistical model to calculate the neutron-capture cross sections and therefore the reaction rates. This model is based on the fundamental Bohr hypothesis that the capture process occurs by means of the intermediary production of a compound system that can reach a state of thermodynamic equilibrium. Provided the reaction $A+n\rightarrow B^{x}+\gamma$ represents the neutron capture onto the nucleus $A$ leaving the residual nucleus $B$ in the $x$th excited state and the photon, the corresponding cross section can be expressed as [20]
$\sigma_{A+n\rightarrow B^{x}+\gamma}(E)=$ $\frac{\pi}{k^{2}}\sum^{l_{max}+I_{A}+I_{n}}_{J=mod(I_{A}+I_{n},1)}\sum^{1}_{\Pi=-1}\frac{2J+1}{(2I_{A}+1)(2I_{n}+1)}$
×$\times\sum^{J+I_{A}}_{J_{n}=\left|J-I_{A}\right|}\sum^{J_{n}+I_{n}}_{l_{i}=\left|J_{n}-I_{n}\right|}\sum^{J+I^{x}_{B}}_{\lambda=\left|J-I^{x}_{B}\right|}\sum^{\lambda+I_{\gamma}}_{l_{f}=\left|\lambda-I_{\gamma}\right|}\delta^{\pi}_{C_{n}}\delta^{\pi}_{C_{\gamma}}$
×$\times\frac{\left\langle T^{J}_{C_{n},l_{i},J_{n}}(E)\right\rangle\left\langle T^{J}_{C_{\gamma},l_{f},\lambda}(E_{\gamma})\right\rangle}{\sum_{Clj}\delta^{\pi}_{C}\left\langle T^{J}_{C,l,j}(E_{C})\right\rangle}W^{J}_{C_{n}l_{i}J_{n}C_{\gamma}l_{f}\lambda}.$ (1)
In the above equations, $E$ ($E_{\gamma}$), $I_{n}$ ($I_{\gamma}$), $\pi_{n}$ ($\pi_{\gamma}$), $l_{i}$ ($l_{f}$), and $J_{n}$ ($\lambda$) represent the energy, spin, parity, orbital, and total angular momentum for the projectile neutron (emitted photon), respectively, and $l_{max}$ is the maximum value of $l_{i}$. The symbols $I_{A}$, $\pi_{A}$ ($I_{B}^{x}$, $\pi_{B}^{x}$) represent the spin and parity of the target (residual) nucleus, while $J$ and $\Pi$ correspond to the spin and parity of the compound system. The initial system ($A+n$) is designated by the channel label $C_{n}=(n,I_{n},E,E_{A}=0,I_{A},\pi_{A})$, and $C_{\gamma}=(\gamma,I_{\gamma},E_{\gamma},E_{B}^{x},I_{B}^{x},\pi_{B}^{x})$ is similar for the final system ($B^{x}+\gamma$), where $E_{A}$ and $E_{B}^{x}$ are the excited energy for target and residual nucleus, respectively. $\delta^{\pi}_{C_{n}}=1$ if $\pi_{A}\pi_{n}(-1)^{l_{i}}=\Pi$ and 0 otherwise. $\delta^{\pi}_{C_{\gamma}}=1$ if $\pi^{x}_{B}\pi_{\gamma}(-1)^{l_{f}}=\Pi$ and 0 otherwise. $k$ is the wave number of the relative motion. $W$ is the width fluctuation correction factor. $T$ is the transmission coefficient and ${\sum_{Clj}\delta^{\pi}_{C}\left\langle T^{J}_{C,l,j}(E_{C})\right\rangle}$ is the sum of transmission coefficient $T$ for all possible decay channels.
Each transmission coefficient $T$ is calculated for all levels with known energy, spin, and parity. If the excitation energy $E^{x}$, which is implicit in the definition of the channel $C$, corresponds to a state in the continuum, an effective transmission coefficient for an excitation-energy bin of width $\Delta E$ is defined by the integral
[…]
over the nuclear level density $\rho(E,J,\Pi)$. $T$ relies on the optical potentials of the two reacting particles. In this work, we use the WLH optical potential from Ref. [52] to calculate $T$ for the neutron. The parameters of the WLH potential are sampled within their respective uncertainty ranges, and the corresponding reaction observables are computed for each sample. The resulting uncertainties in the calculated observables are quantified through extensive sampling and computations. The calculations are performed with the reaction code TALYS [55].
3 Calculations and results
3.1 Comparison with experimental result
In order to test the capability of the WLH potential in describing the cross sections of neutron-capture reactions, we compare in Fig. 1 our calculated results, obtained by sampling the potential parameter set 400 times, with experimental data for the nuclei ${}^{61}\text{Ni}$, ${}^{122}\text{Sn}$, and ${}^{176}\text{Lu}$. A rather good agreement is found, particularly in the low-energy region, with the experimental data primarily lying within the calculated uncertainty ranges, which is especially significant for astrophysical nucleosynthesis.
Figure 1: Comparison of the calculated and experimental neutron-capture cross sections as a function of incident energy $E_{n}$ for three nuclei. The shaded area represents cross sections calculated from 400 random samples of the WLH potential. Experimental data are shown as black points and are taken from Refs. [56] for ${}^{61}\text{Ni}(\text{n},\gamma)^{62}\text{Ni}$, [57, 58] for ${}^{122}\text{Sn}(\text{n},\gamma)^{123}\text{Sn}$, and [59, 60, 61] for ${}^{176}\text{Lu}(\text{n},\gamma)^{177}\text{Lu}$.
Figure 2: Ratios of theoretical to experimental ($n,\gamma$) MACS for the 221 ($26\leq Z\leq 83$) known KADoNiS values at 30 keV. Black squares are calculated using the phenomenological KD potential. Red circles represent the results with the WLH potential, with error bars indicating the maximum and minimum values under the constraint $f_{rms}\leq 1.56$. The gray shaded area guides the eye for deviations within a factor of two.
To further assess the global predictive ability of the WLH potential against all available experimental data, we calculate the Maxwellian-averaged cross sections (MACS) at 30 keV for 221 nuclei with $26\leq Z\leq 83$, and compare the results with experimental data taken from the KADoNiS database [62]. The average values of the WLH potential parameter set are used in the calculations. Other model parameters, such as those for level density and photon strength functions, are taken, wherever possible, from the recommended values in Ref. [63], which are suggested for TALYS calculations in astrophysics applications. As shown by the red circles in Fig. 2, the deviations between the MACS values calculated using the WLH potential and those reported in KADoNiS are generally within a factor of two. Furthermore, we present the theoretical results obtained using the phenomenological potential of Koning and Delaroche (KD) which is widely employed in nuclear reaction calculations and is well known for its good agreement with experimental data through parameter adjustment [38]. It can be seen in Fig. 2 that the results based on the WLH potential are consistent with those obtained from the KD potential in both magnitude and overall trend. These observations confirm the capability of the WLH potential to reproduce experimental data for stable nuclei and support its application to to unstable neutron-rich nuclei.
Sampling of the potential parameter set introduces an uncertainty range in the calculated results. To quantify the predictive accuracy of the WLH potential for each sample, we adopt the root-mean-square (rms) criterion, i.e. $f_{rms}$ deviation, defined in Ref. [63, 64] as
$f_{rms}=\exp\left[\frac{1}{N_{e}}\sum_{i=1}^{N_{e}}\ln^{2}\left(\frac{\left\langle\sigma_{\mathrm{th},i}\right\rangle}{\left\langle\sigma_{\exp,i}\right\rangle}\right)\right]^{1/2},$ (3)
where $N_{e}$ is the number of known MACS from KADoNiS database, and $\left\langle\sigma_{\mathrm{th},i}\right\rangle$ and $\left\langle\sigma_{\exp,i}\right\rangle$ are the theoretical and experimentally MACS, respectively. With Eq. (3), each sampling corresponds to an $f_{rms}$ deviation value, which is always greater than or equal to 1. The deviation from 1 indicates, on average, how far the prediction deviates from the experimental data. In Fig. 3, the distribution of $f_{rms}$ deviations with respect to the 221 experimental ($n,\gamma$) MACS at 30 keV from the KADoNiS database is displayed. As can be seen, while the $f_{rms}$ deviations reach up to 2.16, most are below 1.74. These results are consistent with those reported in Refs. [63, 64], which are based on uncertainty analyses of neutron-capture reactions relevant to astrophysics. The fact that the uncertainties remain within a reasonable range further confirms the validity of the WLH potential in describing neutron-capture experimental data.
Figure 3: Distribution of $f_{rms}$ deviations with respect to the 221 experimental ($n,\gamma$) MACS for 400 samples of the parameter set.
3.2 Application to neutron-rich nuclei
Applying the WLH potential to neutron-capture reactions of nuclei far from stability, for which no experimental data are available, leads to increasingly intensive computations. This is due to the large number of neutron-rich nuclei involved, each requiring extensive sampling of potential parameters across a wide uncertainty range, along with the corresponding reaction calculations. For the purposes of efficiently propagating parameter uncertainties to calculated observables, as well as reducing computational cost, we adopt the backward-forward Monte Carlo (BFMC) method [64, 65, 66, 67]. This method relies on the sampling of the model parameters and the use of a generalized $\chi^{2}$ estimator to quantify the likelihood of each simulation with respect to a given set of experimental constraints. It consists primarily of two Monte Carlo (MC) steps: in the backward MC step, suitable parameter samples are selected based on experimental constraints; in the forward MC step, the selected parameter samples are used to calculate unmeasured quantities.
We use the $f_{rms}$ deviation as the $\chi^{2}$ estimator with respect to the 221 experimental MACS values in the backward MC step. Specifically, we apply a threshold function
$w=\left\{\begin{aligned} &1\quad\mathrm{for}\quad f_{rms}\leq f_{rms}^{crit},\\ &0\quad\mathrm{for}\quad f_{rms}>f_{rms}^{crit},\end{aligned}\right.$ (4)
to select the sampled potential parameter sets for which the $f_{rms}$ deviation is less than or equal to the critical value $f_{rms}^{crit}$. This ensures that the corresponding calculated observables are in closer agreement with the experimental data than those with $f_{rms}>f_{rms}^{crit}$. In Fig. 2, the error bars of the WLH potential are derived from potential parameters constrained by $f_{rms}\leq 1.56$. It can be seen that the uncertainty range of the WLH potential effectively encompasses the theoretical results of the KD potential.
Figure 4: The neutron-capture rates, constrained by three $f_{rms}^{crit}$ values (1.50, 1.56 and 1.74) corresponding to three differently shaded areas, as a function of temperature for the isotopes ${}^{122}\text{Sn}$, ${}^{132}\text{Sn}$, and ${}^{150}\text{Sn}$.
In the forward MC step, the parameter sets selected in the backward MC step are used to calculate the neutron-capture rates of neutron-rich nuclei. As shown in Fig. 4, the neutron-capture rates and their uncertainty ranges as a function of temperature for the Sn isotopes ${}^{122}\text{Sn}$, ${}^{132}\text{Sn}$, and ${}^{150}\text{Sn}$ are calculated using three $f_{rms}^{crit}$ values. As $f_{rms}^{crit}$ decreases, the uncertainty range gradually narrows, indicating that the $f_{rms}$ deviation effectively controls the uncertainty propagation through the selection of potential parameters. For ${}^{122}\text{Sn}$ and ${}^{132}\text{Sn}$, the uncertainty ranges are within one order of magnitude. For ${}^{150}\text{Sn}$, however, the differences in the reaction rates reach up to five orders of magnitude. Fig. 5 shows the neutron-capture rates of Sn isotopes as a function of neutron number $N$ at a temperature of $T=1$ GK. It is more clearly seen that the uncertainty in the rates decreases with decreasing $f_{rms}^{crit}$, and when $N$ exceeds approximately 87, the uncertainty increases dramatically and exhibits an odd-even effect.
Figure 5: The neutron-capture rates for Sn isotopes as a function of neutron number $N$ at a temperature of $T=1$ GK. Differently shaded areas represent different $f_{rms}^{crit}$ values.
Furthermore, using the parameter sets constrained by $f_{rms}\leq 1.56$, we comprehensively calculate the neutron-capture rates at $T=1$ GK for all nuclei with $26\leq Z\leq 84$, located between the valley of stability and the neutron drip line. The ratio of maximum to minimum rates is shown in the chart of the nuclides in Fig. 6. There is a noticeable separation around an isospin asymmetry, i.e., $\delta=(N-Z)/A$, of 0.28, below which the uncertainty range in the reaction rate remains within one order of magnitude, and above which it mostly exceeds three orders of magnitude. A significant effect of the isospin asymmetry on the astrophysical neutron-capture reactions, especially for nuclei off stability line, has been already suggested [68] and recently corroborated by experimental evidence [69]. The separation related to isospin asymmetry in Fig. 6 further confirms this effect. Although the WLH potential parameterizes a global optical potential for each chiral force that depends smoothly on the projectile energy as well as the target nucleus mass number and isospin asymmetry, the isospin dependence of this potential may not be adequately accounted for in nuclei far from the valley of stability. Therefore, future studies focusing on the isospin dependence of the WLH potential in nuclei far from stability are suggested.
Figure 6: Representations in the ($N,Z$) plane of the ratio between the maximum and minimum reaction rates for all nuclei with $26\leq Z\leq 84$, lying between the valley of stability and neutron drip line, at $T=1$ GK and $f_{rms}\leq 1.56$.
4 Conclusions
In this work, we apply the microscopic global optical potential of chiral effective field theory (the WLH potential) to the study of astrophysical neutron-capture reaction for the first time. By comparing the neutron-capture cross sections calculated using the WLH potential with experimental data for the nuclei ${}^{61}\text{Ni}$, ${}^{122}\text{Sn}$, and ${}^{176}\text{Lu}$, we find that the theoretical predictions show good agreement with the experimental results. Furthermore, for all experimentally known MACS with $26\leq Z\leq 83$, the WLH potential also provides a satisfactory description. We employ the $f_{rms}$ deviation to assess the predictive accuracy of the WLH potential for each sample of the parameter set, and find that the associated uncertainty lies within a reasonable range. These results demonstrate the capability of the WLH potential to reliably reproduce experimental data and establish a foundation for its application to unstable neutron-rich nuclei.
For neutron-rich nuclei, to efficiently propagate parameter uncertainties to reaction rates and reduce computational cost, we adopt the BFMC method, using the $f_{rms}$ deviation to constrain parameter sampling. With this approach, we calculate the neutron-capture rates for nuclei ${}^{122}\text{Sn}$, ${}^{132}\text{Sn}$, and ${}^{150}\text{Sn}$ at different temperatures, as well as the rates for Sn isotopes as a function of neutron number at a temperature of $T=1$ GK. It is found that the $f_{rms}$ constraint effectively controls the propagation of uncertainty, and a sharp increase in uncertainty is observed around $N=87$. A further comprehensive calculation, covering all nuclei with $26\leq Z\leq 84$ between the valley of stability and the neutron drip line at $T=1$ GK and constrained by $f_{rms}\leq 1.56$, shows that the WLH potential provides a good description with relatively small uncertainties for nuclei with isospin asymmetry $\delta<0.28$. However, for nuclei with $\delta>0.28$, the uncertainty in the reaction rates increases significantly. This highlights the critical role of isospin dependence in optical potentials and suggests that future developments of the WLH potential may pay greater attention to this aspect.
5 Acknowledgements
This work is supported by the National Key Research and Development program (MOST 2022YFA1602304), the National Natural Science Foundation of China (Grants No. 12335009, 12175156, and 12375122), the Strategic Priority Research Program of Chinese Academy of Sciences (Grant No. XDB34020200), and the Fundamental Research Funds for the Central Universities (Grant No. B240201048). Y. X. acknowledges support from the Romanian Ministry of Research, Innovation and Digitalization under Contract No. PN 23 21 01 06 and the ELI-RO project with Contract ELI-RORDI-2024-008 (AMAP), and support from the IAEA CRP on ”Updating and Improving Nuclear Level Densities for Applications (Project No. F41034)”.